A 500 MW steam power plant consumes coal at a rate of 65 tons/h. If the heating value of the coal is 30,000 kJ/kg, what is the overall efficiency of this plant? 


1
Expert's answer
2021-12-02T20:23:00-0500

"Q_{in} =M_{coal} \u00d7q_{hv}"

Where:

"Q_{in} =" Amount of heat input

"M_{coal} =" Mass of coal consumed per day

"q_{hv} =" heating value

"Q_{in} = 65000kg\/hr \u00d724hr\u00d730MJ\/Kg=4.68\u00d710^7 MJ"

Further "Q^.= \frac{Q_{in}}{24 hr}"

Where "Q^.=" Coal consumption rate

"Q^.= \frac{4.68×10^7MJ}{24×3600}=541.67MJ/s"

Efficiency"n=\frac{ W_{output}}{Q^.}"

"n=\frac{500MW}{541.67MJ/s}=0.923"

"n=" 92.3%